OpenStudy (princesssleelee):

Rationalize the denominator of square root of negative 16 over open parentheses 1 plus i close parentheses plus open parentheses 6 plus 3 i. quantity of negative 16 plus 28 i over 33 quantity of 16 plus 28 i over 65 quantity of negative 12 plus 28 i over 40 quantity of 12 plus 28 i over 58

OpenStudy (princesssleelee):

@jhonyy9

jhonyy9 (jhonyy9):

can you rewrite it using numbers ?

jhonyy9 (jhonyy9):

so i will try it too

OpenStudy (princesssleelee):

\[\frac{ \sqrt{-16}}{ (1+i)+(6+3i) }\]

jhonyy9 (jhonyy9):

so you have wrote it right sure ?

jhonyy9 (jhonyy9):

so do you know what mean rationaliz the denominator ?

OpenStudy (princesssleelee):

that is the square root of -16 as the numerator

OpenStudy (princesssleelee):

And the denominator is correct

OpenStudy (princesssleelee):

The reciprocal is created by inverting the numerator and denominator of the starting expression. Since we now have a radical in the denominator, we must rationalize this denominator. Multiply top and bottom by the conjugate of the denominator

OpenStudy (princesssleelee):

Thats from an example on google.

jhonyy9 (jhonyy9):

no - for example : sqrt 3 --------------- = (3-2i) +(1-i ) so rationaliz the denominator mean that you need to do anything for eliminate the terms with ,,i" from denominator - yes ? in this case multiplie numerator and denominator by conjugate of denominator do you know it how and what mean conjugate ?

OpenStudy (princesssleelee):

yes

jhonyy9 (jhonyy9):

so than try it so

OpenStudy (princesssleelee):

idk how

jhonyy9 (jhonyy9):

you have said that you know what mean conjugate - yes ?

OpenStudy (princesssleelee):

i had the incorrect definition

jhonyy9 (jhonyy9):

so suppose there is in denominator (a+ib) + (c+id) - and because you know formula (a+b)(a-b)=a^2 -b^2 - and you know again that i^2 = -1 - so try using these what i have wrote above ok ?

OpenStudy (princesssleelee):

ok

jhonyy9 (jhonyy9):

than what will be the conjugate of denominator in case of your exercise ?

jhonyy9 (jhonyy9):

so there are (1+i)+(6+3i) (a + b) - let being so - and you know formula (a+b)(a-b)=a^2 -b^2 us this please

jhonyy9 (jhonyy9):

so you need multiplie the numerator and the denominator by (a-b) yes ?

jhonyy9 (jhonyy9):

so what will be (a-b) in this case of your exercise ?

jhonyy9 (jhonyy9):

so than you know a=(1+i) and b=(6+3i) - so a-b ho will be ?

jhonyy9 (jhonyy9):

how - sorry

jhonyy9 (jhonyy9):

do you understand it ?

jhonyy9 (jhonyy9):

so the conjugate will be (1+i)-(6+3i) yes ?

jhonyy9 (jhonyy9):

multiplie by this the numerator and denominator - ok. ?

jhonyy9 (jhonyy9):

((1+i)+(6+3i))*((1+i)-(6+3i)) = (1+i)^2 -(6+3i)^2 do you like continue it ? but dont forget it that this is just the denominator and you need again multiplie by conjugate of denominator the numerator too

jhonyy9 (jhonyy9):

@princesssleelee are you here please ?

OpenStudy (princesssleelee):

im here

OpenStudy (radar):

First simplify getting: